# LLM $\TITLELLM for Large-Scale Optimization Model Auto-Formulation$: A Lightweight Few-Shot Learning Approach

### v1

@@ 3.2 Workflow Generation and Model Generation
At a high level, the workflow generation agent picks a demo instance from the Ref-Data that is contextually and structurally similar to the user’s input. This is then structured into several components to form the workflow. In this section, we present two kinds of workflows, one type-tailored and one type-agnostic, to accommodate different scenarios. In particular, we use type-tailored workflow for all the problems except for those classified as ”Others” and “Mixture”, and use type-agnostic workflow for these two.
@@ 3.2.1 Type-Tailored Workflow
  • Action and Action Input: This step invokes the use of the customized RAG tool CSVQA and is a crucial component in data-retrieval. With this tool, the input datasets are first converted into vector representations line by line. When a query comes in, it applies similarity search to find lines of data that best match the needs of the problem description. For instance, the demo example $q$ asks for information related to Sony products (see Section E of the appendix). CSVQA thus retrieves all the products related to the keyword “Sony” while excluding irrelevant information (e.g., “Apple” product, which is also brought up in the context, but not used in the final model). Notably, CSVQA supports fuzzy search, which is often required in optimization modeling;
@@ 3.3 Novelties in Our Design and Key Messages
The Value of Agentic Workflow Construction: As discussed earlier, a central challenge in applying LLMs to large-scale optimization model auto-formulation is reasoning over long, heterogeneous, and often unstructured token sequences. To mitigate this challenge, LEAN-LLM-OPT designs workflows that decompose a complex optimization modeling task into a sequence of structured sub-tasks, facilitating partial standardization. Moreover, given the versatility of optimization modeling, the workflows are constructed dynamically and leverage LLMs’ natural language processing capabilities. Specifically,
@@
  • Leveraging LLMs’ language understanding capabilities, both workflows are instantiated around a reference problem, helping the downstream agent identify which parts of the datasets merit closer examination and form an initial hypothesis of the target optimization model, rather than reasoning entirely from scratch.
As a result, the workflows help direct the model-generation agent’s attention to the most challenging components of the modeling task (e.g., interpreting the query intent, aligning coefficients across the objective and constraints, and incorporating subtle but critical constraints), enabling the optimization model to be constructed more incrementally and effectively.
@@
Despite these gaps, our framework can still utilize Ref-Data effectively, suggesting potential for extension to a broader range of applications. Moreover, our plug-and-play strategy allows LEAN-LLM-OPT to be readily paired with more advanced LLMs as they become available.
@@ 4 Numerical Simulations
Metrics and Solver Configurations: For fair comparisons, we follow the convention in Huang et al. (2025) to use the execution accuracy as our evaluation metric. The execution accuracy is defined as the ratio between the number of an executed optimal value that matches the provided ground truth optimal value and the total number of problems [^4] See Remark 2 for more details of this automated process for LEAN-LLM-OPT..
@@
Table 5: Overall performance and performance by problem type. LEAN is the abbreviation of LEAN-LLM-OPT. The best performing model in each column is boldfaced and the second best is underlined
@@
ORLM 32% 68.18% 33.33% 14.29% 60% 12.5% 5.56% 32.67%
OptiMUS 0% 0% 0% 21.43% 60% 0% 5.56% 6.93%
Gemini 3 Pro 12% 63.64% 100% 50% 80% 75% 55.56% 52.48%
GPT-5.2 4% 68.18% 77.78% 35.71% 40% 50% 38.89% 40.59%
GPT-5 40% 13.64% 33.33% 35.71% 80% 12.50% 27.78% 30.69%
GPT-4.1 20% 18.18% 22.22% 21.43% 0% 0% 5.56% 14.85%
gpt-oss-20B 0% 45.45% 0% 28.57% 40% 25% 44.44% 25.74%
LEAN (GPT-4.1) 72% 86.36% 100% 85.71% 100% 100% 83.33% 85.15%
LEAN (gpt-oss-20B) 80% 90.91% 100% 64.29% 60% 75% 77.78% 80.2%
LEAN (GPT-4.1-Agnostic) 20% 68.18% 77.78% 7.15% 80% 100% 83.33% 62.35%
LEAN (gpt-oss-Agnostic) 12% 59.09% 88.89% 21.43% 60% 75% 77.78% 56.31%
@@
Results on Large-Scale-OR: Table 5 summarizes the execution accuracy of different models on Large-Scale-OR (LEAN is the abbreviation of LEAN-LLM-OPT). We report the overall results as well as the detailed results on different problem types. Both LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) achieve overall accuracies of above 80%, significantly exceeding the competing methods’ execution accuracies. It is worth mentioning that their base models, GPT-4.1 and gpt-oss-20B, could only achieve overall accuracies of less than 26%. This means LEAN-LLM-OPT is able to dramatically improve their performance by stimulating their reasoning capabilities with the workflows.
@@
Among the competing methods, Gemini 3 Pro has the best execution accuracy of 52.48%. The fine-tuned model ORLM has an overall execution accuracy of 32.67% while OptiMUS is only able to achieve a less than 7% execution accuracy.
@@
Recall that in LEAN-LLM-OPT, only problems classified as “Others” and “Mixture” would be handled by the type-agnostic workflow. To perform a robustness check, we manually route all problems to the type-agnostic workflow. The results are shown under LEAN-LLM-OPT (GPT-4.1-Agnostic) and LEAN-LLM-OPT (gpt-oss-Agnostic). As we can see, they perform worse than LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B), but still manage to outperform all other competing methods. These showcase the effectiveness of our type-tailored workflow and also the robustness of our type-agnostic workflow.
@@
By inspecting the relationship between modeling accuracy and length of input (see Figure 4), we find that one of the key reasons for the excellent performance of LEAN-LLM-OPT comes from its capability to sustain longer inputs. We can also find from Figure 5 that LEAN-LLM-OPT’s performance drops as the number of variables in the output increases. This is expected because as the number of variables increases, LEAN-LLM-OPT has to figure out the placement of more coefficients. But somewhat surprisingly, several methods’ performance can increase as the number of variables of the output increases. This suggests 1) similar to other reasoning tasks, the difficulty in large-scale optimization modeling may also come from the size of the input; 2) there might be better ways to handle the placement of coefficients, which can further improve the performance of LEAN-LLM-OPT as well.
Figure 4: Modeling accuracy vs. input size
Figure 5: Modeling accuracy vs. output size
@@
Table 6: Performance comparison of models on benchmarks. “–” means the model was unable to output a complete and valid result. Results labeled with “*” are acquired from Chen et al. (2025b). The best performing model in each column is boldfaced and the second best is underlined
@@
gpt-oss-20B 35.1% 67.9% 27.4% 19.0%
LEAN-LLM-OPT (GPT-4.1) 94.7% 93.9% 60.1% 56.0%
LEAN-LLM-OPT (gpt-oss-20B) 93.5% 91.4% 40.4% 59.0%
Results on Small-Scale Benchmarks: Table 6 summarizes the results on the small-scale benchmarks. Comparing LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) against the base models, we can again see that LEAN-LLM-OPT can bring substantial improvements in modeling capabilities. In terms of absolute modeling accuracies, LEAN-LLM-OPT (GPT-4.1) performs the best in NL4OPT and MAMO Easy; while Gemini 3 Pro performs the best in MAMO Complex and IndustryOR. LEAN-LLM-OPT (gpt-oss-20B) achieves the second best in NL4OPT and IndustryOR and GPT-5.2 achieves the second best in MAMO Easy and MAMO Complex. Compared with ORLM and OptiMUS, LEAN-LLM-OPT (GPT-4.1) achieves higher performance across all benchmarks, while LEAN-LLM-OPT (gpt-oss-20B) trails OptiMUS on the MAMO Complex benchmark.Through these results, LEAN-LLM-OPT demonstrate its ability to handle small-scale optimization modeling tasks.
@@ 5 Case Study: Singapore Airlines Choice-Based Revenue Management
In this section, we demonstrate the practicality of LEAN-LLM-OPT using a case study on Singapore Airlines’ choice-based revenue management and we consider a generalized SBLP formulation (Gallego et al. 2015). We are interested in examining LEAN-LLM-OPT’s ability to earn revenue in this setting. In this case study, our model faces a more challenging task because it has to handle multiple datasets and the target model is more complicated than previous ones. In this section, we include Gemini 3 Pro and GPT-5.2 as our competitors because they are the best performing ones in the previous rounds of simulations. For fair comparison, our competing methods are provided with one example that includes the problem description, the correct formulation, the Python code, and the original research paper by Gallego et al. (2015).
@@ 5.2 Available Data and Testing Dataset Air-NRM
In Air-NRM-CA, one is asked to formulate the SBLP (2) that optimizes fare type capacity for all flights chosen by the query (see Section H for an example query). In Air-NRM-NP, it is further required that the optimization formulation should pick a subset of flights (e.g., 3 flights among all those available) that maximize revenue and ensure the flow conservation constraint holds. Below is an example query in Air-NRM-NP,
@@ 5.3 Performance on Fare Type Capacity Allocation and Ablation Study
Table 11: Execution accuracy of different approaches on Air-NRM-CA
Model Execution accuracy
Gemini 3 Pro 80.00%
GPT-5.2 26.67%
LEAN-LLM-OPT (GPT-4.1) 93.33%
LEAN-LLM-OPT (gpt-oss-20B) 86.67%
LEAN-LLM-OPT (GPT-4.1, tools-only) 20.00%
LEAN-LLM-OPT (GPT-4.1, workflow-only) 0.00%
LEAN-LLM-OPT (GPT-4.1, base model-only) 0.00%
In this section, we report the performance of LEAN-LLM-OPT on the fare type capacity allocation dataset Air-NRM-CA. Table 11 presents the execution accuracy of various approaches. In this table, LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) exhibit leading performance compared to the other methods. Between the competing methods, we can also see that Gemini 3 Pro has a much stronger performance when compared to GPT-5.2.
@@
The tools-only approach has an execution accuracy of 20.00%. With access to the data-handling tools, it can retrieve relevant data, such as flight capacities, average passenger counts, ticket prices, and attraction values from the input datasets. However, this approach often struggles with interpreting the retrieved data and placing the coefficients correctly, especially when the problem size grows larger. We reason that this is because the model generation agent is not provided with a step-by-step guide on data usage. Note that in this use case, the query comes with multiple datasets; it is thus important for the agent to correctly interpret the retrieved data so that it knows which equation every piece of data belongs to.
The execution accuracy of the workflow-only approach is 0.00%. Although this approach can adeptly interpret the retrieved data, a critical drawback is that it can hardly extract all the required data even when the size of the problem instance is small. This is because it could only use Python built-in functions rather than the customized tools. This drastically degrades the capability in data handling. Therefore, it can only correctly generate the abstract model without correct and complete parameters.

### v2

@@ 3.2 Workflow Generation and Model Generation
At a high level, the workflow generation agent picks a demo instance from the Ref-Data that is contextually and structurally similar to the user’s input. This is then structured into several components to form the workflow. In this section, we present two kinds of workflows, one type-tailored and one type-agnostic, to accommodate different scenarios. In particular, we use type-tailored workflow for all the problems except for those classified as “Others” and “Mixture”, and use type-agnostic workflow for these two.
@@ 3.2.1 Type-Tailored Workflow
  • Action and Action Input: This step invokes the use of the customized RAG tool CSVQA and is a crucial component in data-retrieval [^4] We note that the tool can be different depending on our base model and the type of the problem.. With this tool, the input datasets are first converted into vector representations line by line. When a query comes in, it applies similarity search to find lines of data that best match the needs of the problem description. For instance, the demo example $q$ asks for information related to Sony products (see Section E of the appendix). CSVQA thus retrieves all the products related to the keyword “Sony” while excluding irrelevant information (e.g., “Apple” product, which is also brought up in the context, but not used in the final model). Notably, CSVQA supports fuzzy search, which is often required in optimization modeling;
@@ 3.2.3 Model Generation Agent
Remark 2 (The Benefit of Classification and Our Design).
We note that the classification agent is not strictly required as LEAN-LLM-OPT can perform well even without it (see the numerical experiment results in Table 6). However, using a classification agent can help customize the subsequent workflow for commonly encountered problem types, and hence, freeing up reasoning bandwidth for the downstream agents to further improve performance. Two of our design choices worth further discussions.
  • First, one may consider directly using RAG without LLM reasoning. However, as we show in Section D of the appendix, the classification accuracy can be negatively impacted if so. This underscores the importance of leveraging the reasoning capability of the LLMs;
  • Moreover, given the large input token limits of modern LLMs, one might argue that we could simply include the entirety of Ref-Data and allow the agents to decide autonomously which parts are relevant. However, as noted in Challenge 1 of Section 1, the performance of state-of-the-art LLMs can degrade substantially once the input length exceeds roughly 800 tokens. Since Ref-Data contains more than 40,000 tokens, directly placing it all into the prompt would likely exacerbate this issue and could further hinder the model’s ability to generate accurate formulations.
@@ 3.3 Novelties in Our Design and Key Messages
Bridging Flexibility and Standardization: As discussed earlier, a central challenge in applying LLMs to large-scale optimization model auto-formulation is reasoning over long, heterogeneous, and often unstructured token sequences, given that it is infeasible to collect sufficient data to cover all possible cases. To mitigate this challenge, LEAN-LLM-OPT follows common practices to design workflows that decompose a complex optimization modeling task into a sequence of structured sub-tasks, facilitating partial standardization. Moreover, given the versatility of optimization modeling, the workflows are constructed dynamically and leverage LLMs’ natural language processing capabilities. Specifically,
@@
  • Leveraging LLMs’ flexible language understanding capabilities, both workflows are instantiated around a reference problem, helping the downstream agent identify which parts of the datasets merit closer examination and form an initial hypothesis of the target optimization model, rather than reasoning entirely from scratch. LLMs’ flexibility also helps to handle the variability in the inputs (e.g., different column names of similar problems), and minimizes the needs for human intervention.
By striking such balance between standardization and flexibility, the workflows help direct the model-generation agent’s attention to the most challenging components of the modeling task (e.g., interpreting the query intent, aligning coefficients across the objective and constraints, and incorporating subtle but critical constraints), enabling the optimization model to be constructed more incrementally and effectively.
@@
Despite these gaps, our framework can still utilize Ref-Data effectively, suggesting potential for generalization as detailed below.
  • By including different problem instances in Ref-Data, one can apply LEAN-LLM-OPT to domains and scenarios that are completely different than the ones would be considered in the current work;
  • Moreover, our plug-and-play strategy allows LEAN-LLM-OPT to be readily paired with more advanced LLMs as they become available. This can make sure we can always leverage the continuously evolving reasoning power of the latest LLMs.
@@ 4 Numerical Simulations
Metrics and Solver Configurations: For all benchmarks, we first follow the convention in Huang et al. (2025) to use the computed optimal value of the output model as our evaluation metric. In particular, the modeling accuracy for a benchmark is defined as
[…]
@@
For Large-Scale-OR, we additionally conduct a manual check on the generated formulations, and compare each of them against the ground truth formulation. This evaluates if the generated formulation is an exact match (EM) to the ground truth except for naming of variables. Formally,
[…]
In our results on Large-Scale-OR, we shall show that modeling accuracy is indeed a good proxy for modeling accuracy (EM).
Table 5: Modeling accuracy (EM) by problem type. LEAN is the abbreviation of LEAN-LLM-OPT. The best performing model in each column is boldfaced and the second best is underlined
@@
ORLM 28% 45.5% 0% 0% 60% 0% 5.6% 19.8%
OptiMUS 0% 18.2% 0% 0% 0% 12.5% 0% 5.0%
Gemini 3 Pro 4% 63.6% 100% 50% 80% 75% 55.6% 50.5%
GPT-5.2 0% 68.2% 77.8% 35.7% 40% 50% 33.3% 38.6%
GPT-5 40% 13.6% 33.3% 35.7% 80% 12.5% 27.8% 30.7%
GPT-4.1 20% 18.2% 22.2% 21.4% 0% 0% 5.6% 14.9%
gpt-oss-20B 0% 45.5% 0% 28.6% 40% 25% 44.4% 25.7%
LEAN (GPT-4.1) 72% 86.4% 100% 78.6% 100% 100% 77.8% 83.2%
LEAN (gpt-oss-20B) 80% 86.4% 88.9% 64.3% 60% 75% 72.2% 77.2%
LEAN (GPT-4.1-Agnostic) 20% 45.5% 44.4% 7.1% 80% 100% 77.8% 45.5%
LEAN (gpt-oss-Agnostic) 12% 54.6% 88.9% 28.6% 40% 75% 72.2% 47.5%
Table 6: Modeling accuracy by problem type. LEAN is the abbreviation of LEAN-LLM-OPT. The best performing model in each column is boldfaced and the second best is underlined
Model NRM (25) RA (22) TP (9) FLP (14) AP (5) Others (8) Mixture (18) Overall
ORLM 32% 68.2% 33.3% 14.3% 60% 12.5% 5.6% 32.7%
OptiMUS 0% 18.2% 0% 0% 0% 12.5% 0% 5.0%
Gemini 3 Pro 12% 63.6% 100% 50% 80% 75% 55.6% 52.5%
GPT-5.2 4% 68.2% 77.8% 35.7% 40% 50% 38.9% 40.6%
GPT-5 40% 13.6% 33.3% 35.7% 80% 12.5% 27.8% 30.7%
GPT-4.1 20% 18.2% 22.2% 21.4% 0% 0% 5.6% 14.9%
gpt-oss-20B 0% 45.5% 0% 28.6% 40% 25% 44.4% 25.7%
LEAN (GPT-4.1) 72% 86.4% 100% 85.7% 100% 100% 83.3% 85.1%
LEAN (gpt-oss-20B) 80% 90.9% 100% 64.3% 60% 75% 77.8% 80.2%
LEAN (GPT-4.1-Agnostic) 20% 68.2% 77.8% 7.1% 80% 100% 100% 54.5%
LEAN (gpt-oss-Agnostic) 12% 59.1% 88.9% 28.6% 60% 75% 77.8% 50.5%
@@
Results on Large-Scale-OR: Table 5 summarizes the modeling accuracy (EM) of different models on Large-Scale-OR (LEAN is the abbreviation of LEAN-LLM-OPT). We report the overall results as well as the detailed results on different problem types. LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) achieve highest overall accuracies, substantially exceeding the competing methods’ modeling accuracies. It is worth mentioning that their base models, GPT-4.1 and gpt-oss-20B, could only achieve overall accuracies of less than 26%. This means LEAN-LLM-OPT is able to dramatically improve their performance by stimulating their reasoning capabilities with the agentic workflows.
@@
Among the competing methods, Gemini 3 Pro has the best modeling accuracy (EM) of 50.5%. The fine-tuned model ORLM has an overall modeling accuracy (EM) of 19.8% while OptiMUS is only able to achieve a 5% modeling accuracy (EM).
@@
Recall that in LEAN-LLM-OPT, only problems classified as “Others” and “Mixture” would be handled by the type-agnostic workflow. To perform a robustness check, we manually route all problems to the type-agnostic workflow. The results are shown under LEAN-LLM-OPT (GPT-4.1-Agnostic) and LEAN-LLM-OPT (gpt-oss-Agnostic). As we can see, they perform worse than LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B), but still manage to outperform almost all other competing methods. These showcase the effectiveness of our type-tailored workflow and also the robustness of our type-agnostic workflow.
@@
By inspecting the relationship between modeling accuracy (EM) and length of input (see Figure 4), we find that one of the key reasons for the excellent performance of LEAN-LLM-OPT comes from its capability to sustain longer inputs. We can also find from Figure 5 that LEAN-LLM-OPT’s performance drops as the number of variables in the output increases. This is expected because as the number of variables increases, the formulation’s complexity also increases, which leaves LEAN-LLM-OPT a more challenging reasoning task (e.g., figure out the placement of more coefficients). But somewhat surprisingly, several methods’ performance can increase as the number of variables of the output increases. This suggests 1) similar to other reasoning tasks, the difficulty in large-scale optimization modeling may also come from the size of the input; 2) there might be better ways to handle the placement of coefficients, which can further improve the performance of LEAN-LLM-OPT as well.
Figure 4: Modeling accuracy (EM) vs. input size
Figure 5: Modeling accuracy (EM) vs. output size
Table 6 also provides the modeling accuracy based on the resulted optimal values. By comparing against Table 5, we find that it is indeed possible for a method to generate an incorrect formulation that can result in a correct optimal value. We provide several examples in Section H of the appendix. But as we can see, the discrepancy between the two metrics are not so significant for most of the methods considered. Therefore, modeling accuracy can indeed be viewed as a good proxy for modeling accuracy (EM).
@@
Table 7: Performance comparison of models on benchmarks. Results labeled with “*” are acquired from Chen et al. (2025b). The best performing model in each column is boldfaced and the second best is underlined
@@
gpt-oss-20B 50.2% 68.9% 29.1% 42.0%
LEAN-LLM-OPT (GPT-4.1) 94.7% 93.5% 71.4% 65.0%
LEAN-LLM-OPT (gpt-oss-20B) 93.5% 91.4% 56.2% 59.0%
Results on Small-Scale Benchmarks: Table 7 summarizes the results on the small-scale benchmarks. Comparing LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) against the base models, we can again see that LEAN-LLM-OPT can bring substantial improvements in modeling capabilities. In terms of absolute modeling accuracies, LEAN-LLM-OPT (GPT-4.1) performs the best in all the benchmarks; while Gemini 3 Pro makes it a tie in IndustryOR. LEAN-LLM-OPT (gpt-oss-20B) achieves the second best in NL4OPT and IndustryOR and GPT-5.2 achieves the second best in MAMO Easy. Compared with ORLM and OptiMUS, both LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) achieve higher performance across all benchmarks. Through these results, LEAN-LLM-OPT demonstrate its ability to handle small-scale optimization modeling tasks.
@@ 5 Case Study: Singapore Airlines Choice-Based Revenue Management
In this section, we demonstrate the practicality of LEAN-LLM-OPT using a case study on Singapore Airlines’ choice-based revenue management and we consider a generalized SBLP formulation (Gallego et al. 2015). We are interested in examining LEAN-LLM-OPT’s ability to earn revenue in this setting. In this case study, our model faces a more challenging task because it has to handle multiple datasets and the target model is more complicated than previous ones. In this section, we include Gemini 3 Pro and GPT-5.2 as our competitors because they are the best performing ones in the previous rounds of simulations. Because SBLP is a relatively less common model, we provide our competing methods with a single example that includes the problem description, the correct formulation, the corresponding Python code, and the original paper by Gallego et al. (2015).
@@ 5.2 Available Data and Testing Dataset Air-NRM
In Air-NRM-CA, one is asked to formulate the SBLP (4) that optimizes fare type capacity for all flights chosen by the query (see Section I for an example query). In Air-NRM-NP, it is further required that the optimization formulation should pick a subset of flights (e.g., 3 flights among all those available) that maximize revenue and ensure the flow conservation constraint holds. Below is an example query in Air-NRM-NP,
@@ 5.3 Performance on Fare Type Capacity Allocation and Ablation Study
Table 12: Modeling accuracy of different approaches on Air-NRM-CA
Model Modeling accuracy
Gemini 3 Pro 80%
GPT-5.2 26.7%
LEAN-LLM-OPT (GPT-4.1) 93.3%
LEAN-LLM-OPT (gpt-oss-20B) 86.7%
LEAN-LLM-OPT (GPT-4.1, tools-only) 20%
LEAN-LLM-OPT (GPT-4.1, workflow-only) 0%
LEAN-LLM-OPT (GPT-4.1, base model-only) 0%
In this section, we report the performance of LEAN-LLM-OPT on the fare type capacity allocation dataset Air-NRM-CA. Table 12 presents the modeling accuracy of various approaches. In this table, LEAN-LLM-OPT (GPT-4.1) and LEAN-LLM-OPT (gpt-oss-20B) exhibit leading performance compared to the other methods. Between the competing methods, we can also see that Gemini 3 Pro has a much stronger performance when compared to GPT-5.2.
@@
The tools-only approach has an modeling accuracy of 20%. With access to the data-handling tools, it can retrieve relevant data, such as flight capacities, average passenger counts, ticket prices, and attraction values from the input datasets. However, this approach often struggles with interpreting the retrieved data and placing the coefficients correctly, especially when the problem size grows larger. We reason that this is because the model generation agent is not provided with a step-by-step guide on data usage. Note that in this use case, the query comes with multiple datasets; it is thus important for the agent to correctly interpret the retrieved data so that it knows which equation every piece of data belongs to.
The modeling accuracy of the workflow-only approach is 0%. Although this approach can adeptly interpret the retrieved data, a critical drawback is that it can hardly extract all the required data even when the size of the problem instance is small. This is because it could only use Python built-in functions rather than the customized tools. This drastically degrades the capability in data handling. Therefore, it can only correctly generate the abstract model without correct and complete parameters.